Your answer is partaly correct. Try again and partie lepenen es an attractive orce of SON, what i^^ vac urn, two partides have charges ot 91 and q, where q.. +3.7HC. They are separated by a distance of 0.21 m the value of 42, with its sign Number T0.00000663
Answer #1

since it is attractive force , the charges would be opposite in nature

so q2 would be a negative charge

F = kq1q2 / r^2

where k = 9*10^9

q1 , q2 are the charges

r = seperation between the charges

5 = 9*10^9*3.7*10^-6*q2 / 0.21^2

q2 = 6.62*10^-6 C

= -6.62 uC